OpenStudy (anonymous):

6 divided by 4/7 a. 6 1/7 b. 5 4/7 c. 10 1/2

OpenStudy (cwrw238):

invert 4/7 then multiply

OpenStudy (anonymous):

c

OpenStudy (anonymous):

because 42 divided by 4 is 10.5

OpenStudy (anonymous):

It can be calculated if not sure of your answer

OpenStudy (anonymous):

Thanks, I also have one more problem I'm having trouble with if that's fine. Write the answer in simplest form: -2/5(-6/7) a. -12/35 b. 6/7 c. 12/35

OpenStudy (anonymous):

c

OpenStudy (tkhunny):

Very clear... \(\dfrac{6}{4/7} = 6\cdot\dfrac{7}{4} = \dfrac{6}{4}\cdot 7 = \dfrac{3}{2}\cdot \dfrac{7}{1} = \dfrac{3\cdot 7}{2\cdot 1} = \dfrac{21}{2} = 10\;\dfrac{1}{2}\) You go ahead and do a "Very Clear..." version on the second problem. Show us yoru work.

OpenStudy (anonymous):

12/35 @clunchin

OpenStudy (tkhunny):

@anastasiad239 Please read the Code of Conduct and conform to that behavior.

OpenStudy (anonymous):

What did i do wrong ? @tkhunny

OpenStudy (cwrw238):

don't just give the answer - explain how to do the problem or give a hint , etc

OpenStudy (cwrw238):

this is how you help people to improve their math